OpenStudy (anonymous):

2 Nm of work is required to stretch a spring from its natural length 30 cm to 42 cm. How much work is needed to stretch the spring from 35 cm to 40 cm?

OpenStudy (anonymous):

1st : this is a physics problem. Next: k=-f/x or k=2/12=1/6. I'm not sure about the actual answer. I'm trying to decide if it is orthodox to take the work required to stretch to 40 cm, then to subtract the work required to stretch to 35 cm.

OpenStudy (anonymous):

That actually just gets me -5/6, which doesn't make sense.

OpenStudy (anonymous):

This is an application problem of integrals. The force used to stretch a spring is from hookes law; what I did: find the force to stretch 30 to 42 from this force find the k (spring constant) with k known, find force to stretch 35 to 40 with this force, find the work but my answer didnt match with the books!

OpenStudy (anonymous):

Try \[W= \int\limits_{35}^{40} -1/6 x dx\]

OpenStudy (anonymous):

which is -125/4